Determine the time it would take to change the temperature by 10°C.

The knowns for conductive heat transfer are given below. First, calculate the rate of conductive heat transfer. Then, determine the time it would take to change the temperature by 10°C.

k = 0.8

A = 10 m^2

T1 = 290 K

T2 = 260 K

d = 0.2 m
Answer #1

The conductive heat transfer is given by

q/t = -kA(dt/dx)

Here, q/t = .8×10×((290-260)/0.2)

q/t = 1200 Watts

To determinete the time required, we need to know the specific heat of the object ( or atleast what the object is ,( whether, air or water or any other)) and also the mass.

The equation of heat is given by

Q = m × Cp × dT

It is only given that dT = 10 degree Celsius

Assuming mass of 1kg and the matter to be water (Cp = 4.18 kJ/ kgK)

Q = 1 × 4180 × 10

Q = 41800 Joules

  

Now the time taken for the increase in temperature is

t = Q ÷ q

= 41800 ÷ 1200

= 34.83 seconds.